# How to Learn in a Noisy World? Self-Correcting the Real-World Data Noise on Machine Translation

### v1

@@ Abstract
The massive amounts of web-mined parallel data contain large amounts of noise. Semantic misalignment, as the primary source of the noise, poses a challenge for training machine translation systems. In this paper, we first study the impact of real-world hard-to-detect misalignment noise by proposing a process to simulate the realistic misalignment controlled by semantic similarity. After quantitatively analyzing the impact of simulated misalignment on machine translation, we show the limited effectiveness of widely used pre-filters to improve the translation performance, underscoring the necessity of more fine-grained ways to handle data noise. By observing the increasing reliability of the model’s self-knowledge for distinguishing misaligned and clean data at the token-level, we propose a self-correction approach which leverages the model’s prediction distribution to revise the training supervision from the ground-truth data over training time. Through comprehensive experiments, we show that our self-correction method not only improves translation performance in the presence of simulated misalignment noise but also proves effective for real-world noisy web-mined datasets across eight translation tasks.
@@
Self-Correcting the Real-World Data Noise on Machine Translation
Yan Meng Di Wu Christof Monz Language Technology Lab University of Amsterdam {y.meng, d.wu, c.monz}@uva.nl
@@ 1 Introduction
The success of machine translation (MT) models is mainly due to the availability of large amounts of web-crawled data. However, publicly available web-mined parallel corpora (bitext) such as CCAligned (El-Kishky et al., 2020), WikiMatrix (Schwenk and Douze, 2017), ParaCrawl (Bañón et al., 2020) and NLLB (team et al., 2022) are shown to be noisy (Kreutzer et al., 2022; Ranathunga et al., 2024). The notable performance drop on NMT when training with injected synthetic noise (Khayrallah and Koehn, 2018) or fine-tuning with CCAligned (Lee et al., 2022), indicates the importance of developing noise-robust training when learning from noisy corpus.
@@
Given a noisy training dataset, a common and straightforward approach to mitigate the impact of noisy data is to filter low-quality training samples (Herold et al., 2022; Bane et al., 2022). However, in practice, large amounts of misaligned samples still exist in pre-filtered web-mined datasets (Kreutzer et al., 2022). Even earlier works (Khayrallah and Koehn, 2018; Herold et al., 2022; Li et al., 2024) study the impact of misalignment on machine translation, but their way of generating synthetic misaligned sentences is by random shuffling, which is unrealistic. The real-world misaligned sentences contain partially shared meanings, which disguise them as seemingly parallel data and are hard-to-detect. To analyze such hard-to-detect real-world misalignment, we design a process to simulate it controlled by semantic similarity. Similar to the real-world setting, we demonstrate that our simulated misalignment also challenges widely-used pre-filters, including Laser (Artetxe and Schwenk, 2018), Comet (Rei et al., 2020) and Bi-Cleaner (Zaragoza-Bernabeu et al., 2022).
@@
Under our simulated noise setting, we evaluate an alternative type of approach to handle data noise: Data truncation (Kang and Hashimoto, 2020; Li et al., 2024; Flores and Cohan, 2024), which focuses on the model training dynamics and ignores losses at the token level when there is a relatively large inconsistency between the model’s prediction and the ground truth during training. However, we observe that truncation methods are sensitive to varying levels of simulated misalignment noise. For example, in low-resource training corpus with a high misalignment rate, truncation methods degrade the translation performance (shown in Section 5.3). We argue that the noisy low-resource setting restricts the model from acquiring sufficient correct knowledge, resulting in an inaccurate removal of clean and useful ground-truth data. Moreover, truncation methods start to ignore potential data noise from an early training time, which overlooks the increasing reliability of the model’s prediction over time.
@@
To overcome these limitations, we propose an approach called self-correction, which leverages the model’s predictions to self-correct noise during training while maintaining supervision from the ground truth to avoid discarding useful training information (as shown in Figure 1). To adapt to the changing reliability of model’s self-knowledge, we set a dynamic schedule to gradually increase the trust in the model’s predictions to revise the training supervision from the ground truth. In the early stage of training, the model is not well trained, so we place greater trust in the reference data than in the model’s prediction. As the model gains knowledge during training, we use it to revise the ground-truth data progressively.
@@
We evaluate our self-correction method in both simulated and real-world environments. We demonstrate that our method consistently outperforms baselines in both high- and low-resource datasets with different levels of misalignment noise. Moreover, we clearly show the gains are mainly from revising the misaligned samples while maintaining the performance of clean parallel data. In the real-world setting, our self-correction method effectively handles naturally occurring noise in web-mined parallel datasets, e.g. ParaCrawl and CCAligned, leading to performance gains of up to 1.2 BLEU points across eight translation tasks and outperforming other alternative methods, including pre-filters and truncation.
@@ 2.1 The Noisy World
Web-mined parallel corpora are the primary training data source for machine translation models. However, parallel data crawled from public websites lack quality guarantees and contain different types of noise (Kreutzer et al., 2022), including wrong language, non-linguistic content, and misalignment noise.
The primary source of noise in parallel web-mined data is semantic misalignment (Khayrallah and Koehn, 2018; Kreutzer et al., 2022; Ranathunga et al., 2024). For instance, Khayrallah and Koehn (2018) analyzed the data quality of the raw ParaCrawl corpus, showing 77% of the analyzed sentence pairs contain noise while half of them are misalignments. Wrong language and non-linguistic contents only account for a small portion since filters e.g., language identification toolkits (Herold et al., 2022), can easily detect them. Kreutzer et al. (2022) extended the data quality analysis to pre-filtered web-mined datasets, e.g. WikiMatrix, CCAligned, noting that more than 50% of data in both corpora are noisy while misalignment is the primary noise. Even in the top-quality parallel data from the NLLB corpus, many misaligned sentences can still be found (Ranathunga et al., 2024).
@@ 2.2 Learning in the Noisy World
2.2.1 Pre-Filter
Data filtering is a straightforward way to remove the noise from the translation training corpus to mitigate its negative impact. There are two methods to ensure the semantic alignments between the source and target in a sentence pair: (1) surface-level filters, e.g., removing sentence pairs that differ a lot in source and target length; (2) semantic-level filters, relying on quality estimation models to score each sentence pair (Kepler et al., 2019; Rei et al., 2020; Peter et al., 2023). Other studies consider the denoising as outlier detection (Taghipour et al., 2011) or ranking problem (Cui et al., 2013).
In this paper, we mainly consider semantic-level filters: (1) Laser (Artetxe and Schwenk, 2018), a sentence alignment pre-filter tool for web-mined corpora, i.e., CCAligned and WikiMatrix; (2) Bi-cleaner (Zaragoza-Bernabeu et al., 2022), a classifier that predicts if a sentence is a translation of another. It is also a part of the pre-processing pipeline for generating the ParaCrawl dataset; (3) Comet (Rei et al., 2020), a widely-used quality estimation model for machine translation.
All these approaches rely on external models to select high-quality data before training. However, our approach does not require such pre-selection but focuses on model training dynamics, which can be applied more broadly.
2.2.2 Truncation
There are two ways for removing data noise by focusing on the model’s training dynamics: (1) using extrinsic trusted data to measure the data quality between noisy and denoised NMT models (Wang et al., 2018); (2) solely using the model’s self-knowledge to remove potential noise during training (Kang and Hashimoto, 2020; Li et al., 2024). In our approach, we do not require any external trusted data or models and thus we only compare with the second type of work, i.e., data truncation.
The most common type of truncation is based on the loss, which estimates the data quality by the model’s predicted probability of the ground truth (Kang and Hashimoto, 2020; Goyal et al., 2022; Flores and Cohan, 2024). Ground-truth tokens with high loss are considered as noise and will be skipped during training by setting their loss to zero. However, relying on loss to remove noise overlooks the model’s probability distribution of non-target tokens. In a high entropy context, the probability of the correct ground-truth token is also low due to many alternative continuations, which cannot be treated as noise. To consider the entire model prediction distribution, Li et al. (2024) further propose Error Norm Truncation and use the norm between the model’s prediction distribution and the one-hot distribution of the ground-truth token to measure the data quality. During training, tokens with error norm values exceeding a pre-defined threshold will be removed.
In this paper, we compare against data truncation methods by relying on two model-based metrics, loss and error norm values. However, truncation methods are sensitive to when and how much data to truncate? Our approach overcomes these limitations by revising all the ground truth with an adaptive trusting weight of the model’s prediction. Details are introduced in Section 4.
@@ 3 An Empirical Study of Misalignment
In this section, we investigate the primary source of noise, i.e., semantic misalignment, in a simulated setting. We first introduce a strategy to simulate realistic misalignment noise by controlling semantic similarity (Section 3.1). Next, we show the similarity of our simulated noise to real-world misalignment in terms of adequacy and its hard-to-detect nature (Section 3.2). Under our simulated noisy setting, we evaluate the model-based metrics to distinguish data noise and show the potential limitations (Section 3.3).
@@ 3.1 Simulating Misalignment Noise
To simulate misalignment, previous works (Bane et al., 2022; Herold et al., 2022; Li et al., 2024) randomly shuffle target sentences of a clean parallel training corpus. However, the random shuffling noise can be easily removed by pre-filters based on length difference or semantic similarity (Herold et al., 2022), which oversimplifies the misalignment in the real-world web-mined corpora. For instance, Kreutzer et al. (2022) show that a large portion of misaligned sentence pairs that share partial semantics still exists even after pre-filter tools, i.e., Laser, (examples in Table 1) which show the challenge of detecting real-world misalignment.
To quantitatively analyze the impact of the realistic misalignment noise, we design a process to simulate real-world misalignment controlled by semantic similarity (See Algorithm 1 in the Appendix). From a chunk of clean data, we first choose misaligned target candidates for a source sentence based on length difference and word overlap from the true parallel target. Next, we use quality estimation methods, e.g., Laser or Comet, to calculate the semantic similarity scores between the source sentence and its misaligned candidates. The candidate with the top score is selected as the final synthetic misaligned target. In this way, the misalignment noise is generated with controlled semantic overlaps. Examples that are controlled by Laser (Misaligned-Laser) and Comet (Misaligned-Comet) are in Appendix 6.
@@ 3.2.1 Adequacy
To show the similarity of our simulated noise to real-world misalignment, we conduct a human evaluation on 50 simulated and real-world misaligned sentences rating their Adequacy (scale 1–5), which measures the meaning overlap between source and target. In Figure 2, we show both the real-world misalignment and simulated misaligned-Laser/Comet has a relatively high adequacy score, above , while Randomly shuffled misaligned sentences only has an adequacy of . This ensures our simulated misalignment contains partial semantic overlaps as the real-world misalignment. Moreover, our simulation process confirms the fluency of misaligned targets by selecting sentences from a clean corpus. Details of human evaluation are in Appendix B.3.
@@ 3.2.2 Hard-to-Detect Nature
To show the hard-to-detect nature of simulated noise as real-world misalignment, we investigate the noise detection ability of widely used pre-filters: Comet, Laser, and Bi-Cleaner.
We calculate the noise detection accuracy of the data filters on a mixed set with the same amounts of clean and noisy data. For the clean data, we randomly sample 2,000 clean sentence pairs from the WMT2017 DeEn training set. For Misaligned-Random, we randomly shuffle the order of target sentences in the sampled clean sentence pairs. For Misaligned-Comet and Misaligned-Laser, we use the same source sentences from the sampled clean data, and select the misaligned targets from another 200K target sentences in the training corpus based on Algorithm 1. We score each sentence pair based on the filter models and determine a true ratio threshold based on the amounts of clean and noisy sentence pairs, here 1:1. Sentence pairs with scores below this threshold are classified as noisy. The details for scoring sentence pairs by filter models are shown in Appendix A.
| Type | Filter Accuracy | ||
| Comet | Laser | Bi-Cleaner | |
| Misaligned-Random | 73% | 76% | 75% |
| Misaligned-Laser | 52% | 40% | 52% |
| Misaligned-Comet | 50% | 60% | 52% |
Table 2 shows the filtering accuracy of data filters on different misaligned noise. First, Misaligned-Random can be accurately detected by all data filters, particularly for using Laser, with an accuracy of 76%. This questions the previous assumptions (Khayrallah and Koehn, 2018; Li et al., 2024) of the impact of misalignment noise on translation performance since most of them can be pre-filtered. However, our introduced noise i.e., Misaligned-Laser and Misaligned-Comet pose a challenge on all pre-filters, as the real-world misalignment does. Overall, we show the validity of our simulated noise in two aspects: (1) Adequacy, reflected in the similar level of shared semantics as the real-world misalignment; (2) Hard-to-Detect Nature, reflected in the low filter accuracy of widely used pre-filters.
@@ 3.3 Fine-grained Misalignment Detection
Model-based metrics, including loss and error norm values, are used in data truncation methods (Kang and Hashimoto, 2020; Li et al., 2024) to detect and ignore data noise at the token-level during training. Here, we evaluate their effectiveness under our simulated noisy settings.
Loss measures the model’s predicted probability of the ground-truth token while error norm value (el2n) calculates the difference between the ground-truth (one-hot) distribution and model’s prediction distribution (eq 1).
Tokens with higher loss or el2n value are indicated as the noise since they are more inconsistent between the model’s prediction and the ground truth.
To evaluate whether the above model-based metrics can distinguish our simulated noise from clean data, we record the loss and el2n values for each token of the 2,000 clean and Misaligned-Laser sentence pairs in the same data setting of Section 3.2.2. Figure 3 shows the changes of loss and el2n value distribution for the clean and noisy tokens in different training epochs. First, we find both metrics can gradually distinguish clean and noisy data as the training time increases. This indicates the effectiveness of the model’s self-knowledge for distinguishing hard-to-detect data noise. Second, we find that the el2n metric shows better distinguishability than loss, indicating the importance of considering the whole model prediction distribution.
However, we point out two limitations of truncation methods relying on model-based metrics: First, they overlook the increasing reliability of model prediction by removing potential data noise from an early training time. Second, they cannot avoid removing clean but useful data. As shown in Figure 3, there is still an undistinguishing area between noisy and clean distribution.

### v2

# How to Learn in a Noisy World? Self-Correcting the Real-World Data Noise in Machine Translation

@@ Abstract
The massive amounts of web-mined parallel data contain large amounts of noise. Semantic misalignment, as the primary source of the noise, poses a challenge for training machine translation systems. In this paper, we first introduce a process for simulating misalignment controlled by semantic similarity, which closely resembles misaligned sentences in real-world web-crawled corpora. Under our simulated misalignment noise settings, we quantitatively analyze its impact on machine translation and demonstrate the limited effectiveness of widely used pre-filters for noise detection. This underscores the necessity of more fine-grained ways to handle hard-to-detect misalignment noise. With an observation of the increasing reliability of the model’s self-knowledge for distinguishing misaligned and clean data at the token level, we propose self-correction—an approach that gradually increases trust in the model’s self-knowledge to correct the training supervision. Comprehensive experiments show that our method significantly improves translation performance both in the presence of simulated misalignment noise and when applied to real-world, noisy web-mined datasets, across a range of translation tasks.
@@
Self-Correcting the Real-World Data Noise in Machine Translation
Yan Meng Di Wu Christof Monz Language Technology Lab University of Amsterdam {y.meng}@uva.nl
@@ 1 Introduction
The success of machine translation (MT) models is mainly due to the availability of large amounts of web-crawled data. However, publicly available web-mined parallel corpora such as CCAligned (El-Kishky et al., 2020), WikiMatrix (Schwenk and Douze, 2017) and ParaCrawl (Bañón et al., 2020) are shown to be noisy (Kreutzer et al., 2022; Ranathunga et al., 2024). The notable performance drop in NMT when training with injected synthetic noise (Khayrallah and Koehn, 2018) or fine-tuning with CCAligned (Lee et al., 2022) indicates the importance of improving the model’s robustness when training with the noisy corpus.
@@
Given a noisy training dataset, a common and straightforward approach to mitigate the impact of noisy data is to filter low-quality training samples (Herold et al., 2022; Bane et al., 2022). However, in practice, large amounts of misalignments still exist in pre-filtered web-mined datasets (Kreutzer et al., 2022). This is because real-world misaligned sentences often share partial meanings, making them appear as seemingly parallel data, increasing the difficulty for pre-filters to detect them. To quantitatively analyze such hard-to-detect real-world misalignments, we design a process to simulate it controlled by semantic similarity. Unlike earlier works (Khayrallah and Koehn, 2018; Herold et al., 2022; Li et al., 2024) that generate misaligned bitext by random shuffling—an approach that is both unrealistic and easy-to-detect—our simulated misalignments closely resemble real-world noise and challenge widely-used pre-filters, such as LASER (Artetxe and Schwenk, 2018) and COMET (Rei et al., 2020).
@@
Under our simulated noise settings, we evaluate a type of approach that could potentially handle misalignment noise: Data truncation (Kang and Hashimoto, 2020; Li et al., 2024; Flores and Cohan, 2024), which ignores losses at the token level during training when there is a relatively large discrepancy between the model’s prediction and the ground truth. Although promising, we observe that truncation methods are sensitive to varying levels of misalignment noise. For example, for low-resource corpora with a high misalignment rate, truncation methods even degrade the translation performance (see Section 5.3). We argue that the noisy low-resource setting prevents the model from acquiring sufficient correct knowledge, resulting in an inaccurate removal of clean and useful ground-truth data. Moreover, truncation methods start to ignore potential data noise from an early training time, which overlooks the increasing reliability of the model’s prediction over time.
@@
To overcome these limitations, we propose an approach called self-correction, which leverages the model’s self-knowledge to correct noise during training while maintaining supervision from the ground truth to avoid discarding useful training information. To adapt to the model’s changing reliability, we set a dynamic schedule to gradually increase the trust in its output. During the early stages of training, we place greater trust in the reference over the model’s predictions. As the model acquires more knowledge, we progressively use the model’s predictions to revise the ground truth.
@@
We evaluate our self-correction method in both simulated and real-world noisy settings. We demonstrate that our method consistently outperforms baselines in both high- and low-resource datasets with different levels of misalignment noise. Moreover, we clearly show the gains are mainly from revising the misaligned samples while maintaining the performance of clean parallel data. In the real-world noise setting, our self-correction method effectively handles naturally occurring noise in web-mined parallel datasets, e.g., ParaCrawl and CCAligned, achieving performance gains of up to 2.1 BLEU points across seven translation tasks and outperforming alternative methods, including pre-filters and truncation.
@@ 2.1 The Noisy World
Web-crawled parallel corpora are the primary training data source for machine translation models. However, parallel data crawled from public websites lack quality guarantees and contain different types of noise (Kreutzer et al., 2022), including wrong language, non-linguistic content, and semantic misalignment.
The primary source of noise in parallel web-mined data is semantic misalignment (Khayrallah and Koehn, 2018; Kreutzer et al., 2022; Ranathunga et al., 2024). For instance, Khayrallah and Koehn (2018) analyzed the data quality of the raw ParaCrawl corpus, showing 77% of the analyzed sentence pairs contain noise with half of them being misalignments. Wrong language and non-linguistic contents only account for a small portion and can be easily handled by filters, e.g., language identification toolkits (Herold et al., 2022). Kreutzer et al. (2022) extended the data quality analysis to pre-filtered web-mined datasets, e.g., WikiMatrix, CCAligned, noting that more than 50% of data in both corpora are noisy while misalignment is the primary noise.
@@ 2.2 Learning in the Noisy World
2.2.1 Data Filter
Data filtering is a straightforward way to mitigate the impact of the noise from translation corpora. Two kinds of filters are often used to ensure the semantic alignment in a sentence pair: (1) surface-level filters, e.g., removing sentence pairs that differ a lot in the source and target length; (2) semantic-level filters, relying on quality estimation models to score each sentence pair (Kepler et al., 2019; Rei et al., 2020; Peter et al., 2023). Other works consider misalignment detection as a ranking problem by training a classifier on annotated synthetic misaligned data (Briakou and Carpuat, 2020).
2.2.2 Training Robustness
The primary limitation of data filters is that they discard entire training samples before training. To retain as much useful information as possible in noisy samples, several methods focus on mitigating their negative impact during model training. For instance, Wang et al. (2018) proposed an online data selection approach that utilizes extrinsic trusted data to identify high-quality samples during training. Similarly, Briakou and Carpuat (2021) employed external semantic divergence tags to guide the training of the translation model. However, both of these approaches depend on external data or factors.
In this paper, we consider an alternative line of works, i.e., data truncation, which relies solely on the model’s self-knowledge to ignore potential noise and further benefits the robustness of model training (Kang and Hashimoto, 2020; Li et al., 2024). For example, Kang and Hashimoto (2020) use losses to estimate data quality, where tokens with high loss are considered as noise and will be ignored during training by setting their loss to zero. Li et al. (2024) further propose Error Norm Truncation, using the norm between the model’s prediction distribution and the one-hot ground-truth token distribution to measure data quality. Their method considers the model’s prediction distribution of non-target tokens, providing a more accurate data quality measurement.
However, there are two limitations of truncation methods: First, they ignore the potential noisy training tokens from a specific training iteration, which overlooks the changes in the model’s reliability during training. Second, ignoring can remove partially clean training information, which can be harmful for low-resource tasks. In this paper, we go a step further and propose a self-correction method to gradually increase the trust of model prediction distributions to correct rather than ignore the ground-truth data during training. Details are introduced in Section 4.
@@ 3 An Empirical Study of Misalignment
In this section, we investigate the primary source of noise, i.e., semantic misalignment, in a simulated setting. We first introduce a strategy to simulate realistic misalignment noise by controlling semantic similarity (Section 3.1). Next, we show the similarity of our simulated noise to real-world misalignment in terms of adequacy and its hard-to-detect nature (Section 3.2). Under our simulated noisy setting, we evaluate model-based metrics to distinguish data noise and highlight their potential limitations (Section 3.3).
@@ 3.1 Simulating Misalignment Noise
To simulate misalignment, previous works (Bane et al., 2022; Herold et al., 2022; Li et al., 2024) randomly shuffle target sentences of a clean parallel corpus. However, random shuffling noise can be easily removed by pre-filters based on length or semantic difference (Herold et al., 2022), which oversimplifies the misalignments found in real-world web-mined corpora. While Briakou and Carpuat (2020) proposed generating fine-grained misaligned targets by perturbing equivalent samples, e.g., deletion or replacement, their method does not guarantee the fluency and authenticity of the misaligned sentences.
To quantitatively analyze the impact of realistic misalignment noise, we designed a process to simulate real-world misalignment controlled by semantic similarity. The main idea is to select misaligned target sentences from a large pool of clean candidates that share partial semantics with the corresponding source sentences, where we use quality estimation models, e.g., LASER or COMET, to measure semantic similarity across languages.
More specifically, given a source sentence and a large pool of target sentences, we first narrow down potential candidates based on the length differences and the word overlap ratio with the true parallel target to reduce computational costs. Then, the candidate with the highest semantic similarity score is selected as the final synthetic misaligned target. By this two-step process, we generate misalignment efficiently while maintaining shared semantics. Algorithm 1 provides a detailed description of our strategy. Examples of misaligned sentences generated using LASER (Misaligned-LASER) and COMET (Misaligned-COMET) can be found in Appendix 6.
@@ 3.2 Real-World Misalignment
| Misaligned Types | Adequacy |
| Real-World | 3.1 |
| Misaligned-COMET | 2.7 |
| Misaligned-LASER | 2.6 |
| Misaligned-Random | 1.2 |
@@ 3.2.1 Adequacy
To show the similarity of our simulated noise to real-world misalignment, we conduct a human evaluation of 200 simulated and real-world misaligned sentences, rating their Adequacy (scale 1–5), which measures the meaning overlap between source and target. In Table 2, we show that both real-world misalignment and Misaligned-LASER/COMET (see Section 3.1) have a relatively high adequacy score, above , while random shuffled misaligned sentences only have an adequacy of . This ensures our simulated misalignment contains only partial semantic overlaps as the real-world misalignment. Details of the human evaluation are in Appendix B.3.
@@ 3.2.2 Hard-to-Detect Nature
To show the hard-to-detect nature of our simulated noise, we investigate the noise detection ability of widely used pre-filters: COMET, LASER, Bi-Cleaner, and XLM-R. The details for each filter model are provided in Appendix A.
We calculate the noise detection accuracy of the data filters on a mixed set with the same amounts of clean and noisy data. For the clean data, we randomly sample 2,000 clean sentence pairs from the WMT2017 DeEn test set. For Misaligned-Random, we randomly shuffle the order of target sentences in the sampled clean sentence pairs. For Misaligned-COMET and Misaligned-LASER, we use the same source sentences from the sampled clean data. We select the misaligned targets from another 200K target sentences in the training corpus based on Algorithm 1. We score each sentence pair based on the filter models and determine a true ratio threshold based on the amounts of clean and noisy sentence pairs, here 1:1. Sentence pairs with scores below this threshold are classified as noisy.
Figure 1 shows the noise detection accuracy of data filters for different misaligned noise. First, all data filters have a relatively high detection accuracy for Misaligned-Random, particularly when using LASER, with an accuracy of 76%. This questions previous assumptions (Khayrallah and Koehn, 2018; Li et al., 2024) of the impact of misalignment noise on translation performance since most of them can be pre-filtered. However, our introduced noise, i.e., Misaligned-LASER and Misaligned-COMET, poses a challenge to all pre-filters, as the real-world misalignment does.
Overall, we show the validity of our simulated noise in two aspects: (1) Adequacy, reflected in the similar level of shared semantics as the real-world misalignment; (2) Hard-to-Detect Nature, reflected in the low noise detection accuracy from widely used pre-filters.
@@ 3.3 Fine-grained Misalignment Detection
To measure the data quality during training, the token-level loss and error norm values are used in data truncation methods (Kang and Hashimoto, 2020; Li et al., 2024). Here, we evaluate their effectiveness under our simulated misalignment noise settings.
Loss measures the model’s predicted probability of the ground-truth token. On the other hand, error norm value (el2n) calculates the difference between the ground-truth (one-hot) distribution and the model’s prediction distribution (eq 1). Tokens with relatively high loss or el2n values are indicated as noise.
We record the loss and el2n values for each token from 2,000 clean and Misaligned-LASER target sentences in the same data setting of Section 3.2.2. Figure 2 shows that clean and misaligned sentences have different loss and el2n distributions as training time increases, from epoch to . This shows the effectiveness of the model’s self-knowledge for distinguishing hard-to-detect misalignment noise from clean sentences.
Notably, the el2n metric exhibits stronger distinguishability compared to loss, underscoring the importance of considering the model’s full prediction distribution. However, the noisy samples’ el2n distribution still partially shifts towards lower values during training, mainly due to the presence of clean tokens in the simulated misaligned sentences. To confirm that the shifted tokens in the noisy samples are truly clean, we provide a token-level annotation (see Appendix B.4) to show that annotated misaligned tokens do have higher el2n values (avg. ) than clean ones (avg. ).
Interestingly, we also observe that clean samples contain tokens with high el2n values (see Table 9). We hypothesize that these tokens might be difficult for the model to learn. Future work could further differentiate between hard-to-learn and noisy tokens and explore their respective impacts on the model’s performance.
Overall, we point out two limitations of truncation methods relying on model-based metrics: First, they overlook the increasing reliability of model prediction by removing potential data noise from an early training time. Second, they cannot avoid ignoring clean but useful data. As mentioned, partial clean tokens still have high el2n values.